# ORACLE: Explaining Feature Interactions in Neural Networks with ANOVA

### v3

@@ 5 Experiments
We compare our ANOVA-based interaction framework, ORACLE, to five SHAP-family (Shapley-based) interaction estimators. ORACLE is not itself a Shapley-based explainer: it does not rely on coalition sampling or game-theoretic axioms, but instead performs an ANOVA-style orthogonal decomposition on a factorial surrogate of the backbone model (Section 3). Nevertheless, ORACLE and the Shapley-based baselines all aim to quantify the same notion of global pairwise interaction strength for each feature pair, which makes them directly comparable in our evaluation. ORACLE fits an ANOVA surrogate on a discretized grid and extracts ANOVA interaction maps and interaction strengths as described in Section 3. As SHAP-based baselines, we include the following methods:
  • FSI: the Faithful Shapley interaction index, which provides an axiomatic Shapley-based measure of global interaction strength (tsai2023faith).
  • SHAP-IQ: a unified Monte Carlo scheme for approximating any-order Shapley interactions (fumagalli2023shap).
  • SVARM-IQ: an accelerated variant of SHAP-IQ that uses stratified sampling over subset sizes to reduce variance and computation (kolpaczki2024svarm).
  • KS-Int: a KernelSHAP-style weighted least-squares formulation for Shapley interaction indices (fumagalli2024kernelshap).
  • DS-Int: a DeepSHAP-style multi-reference interaction approximation built on gradient-based attributions (chen2020explaining).
@@
  • IG@K,B: intervention-oriented utility, defined as the fraction of target ANOVA interaction strengths mass recoverable by selecting up to $B$ pairs among the top-$K$ interactions proposed by a method.
@@ Discussion.
Across the three tabular regression benchmarks, the factorial surrogate consistently recovers the backbone’s top interaction structure more accurately than SHAP-family interaction estimators: ORACLE attains higher NDCG@5 and Xfer-NDCG@5 in nearly all settings, together with strong Peak-IoU@0.10 and CCC. These results indicate better agreement with the target interaction strengths, more stable interaction rankings across independently trained backbones, and well-localized, scale-calibrated detection of interaction hotspots. Beyond these scalar metrics, Figure 2 shows that the recovered main-effect curves match classical DoE-style factor trends, and Figure 3 (Appendix D) reveals localized hotspots aligned with the detected top pairs, suggesting that the recovered low-order structure is visually as well as numerically faithful. The intervention metric IG@K,B is more mixed: SHAP-family methods sometimes achieve similar or slightly higher IG, consistent with their game-theoretic design as marginal-contribution scores tailored to removal-style interventions. In contrast, ORACLE is optimized for $L^{2}$ fidelity of effect surfaces and for capturing low-order structure on the grid, and therefore shows its largest advantages on the synthetic $2^{5}$ factorial design, with still robust gains on Airfoil and kin8nm. Taken together, these findings suggest that moderately complex tabular predictors often admit low-order projections that are faithful enough for interaction benchmarking. As a complementary case study, the latent image and text experiments in Appendix E show that grid-based factorial surrogates are less competitive in highly entangled latent spaces, reinforcing that ORACLE is best suited to settings where features admit interpretable factorial structure.
@@
ORACLE inherits several limitations from its grid-based factorial design. First, interaction maps are defined on a discretized input grid, so performance depends on binning and can degrade when many features or high resolutions lead to sparse cells; our ablation on the grid size $L$ suggests that moderate grids (e.g., $L\in\{3,5,7\}$) offer a reasonable trade-off between resolution and stability. Second, we restrict attention to pairwise interactions in regression backbones and treat the functional-ANOVA decomposition of the trained model under the observed data distribution as the target; higher-order effects, strong feature dependence, or model misspecification can therefore make these targets misaligned with domain notions of interaction. Third, SHAP baselines are evaluated under a fixed Monte Carlo budget and shared reference configuration, and alternative sampling schemes or background choices might improve their metrics and narrow the observed gaps. Finally, our main results focus on low- to medium-dimensional tabular tasks and MLP architectures, so the conclusions may not directly transfer to high-dimensional vision or language models with strong local structure or entangled latent representations.
@@
Several extensions follow naturally from these observations. Methodologically, scaling factorial surrogates beyond low- to medium-dimensional tabular data will likely require adaptive or sparse grids with local refinement around interaction hotspots and structured bases that better respect feature dependence, as well as extensions to classification, sequence, and graph backbones and to higher-order interaction maps. On the evaluation side, combining oracle-based metrics with additional stability and robustness criteria under distribution shift, alternative training runs, or architectural changes could yield more comprehensive interaction benchmarks. Finally, closer integration with SHAP-style estimators—for example, using interaction maps as priors, sanity checks, or guides for targeted Monte Carlo sampling—may lead to hybrid methods that combine the structured semantics of design-of-experiments with the flexibility of Shapley-based interaction indices.
@@ A.1 Additional assumptions and notation
Under Assumptions A.1A.2, the oracle main and pairwise components in the main text can be written as
@@ A.2 Oracle ANOVA decomposition: existence and representation
We sketch the argument. Orthogonality of $H_{0}$ and each $H_{j},H_{jk}$ follows from the centering conditions $\mathbb{E}[h_{j}(X_{j})]=0$ and $\mathbb{E}[h_{jk}(X_{j},X_{k})]=0$. For $H_{j}$ and $H_{\ell}$ with $j\neq\ell$,
$\mathbb{E}\big[h_{j}(X_{j})h_{\ell}(X_{\ell})\big]=\mathbb{E}\big[h_{j}(X_{j})\,\mathbb{E}[h_{\ell}(X_{\ell})\mid X_{j}]\big]=\mathbb{E}\big[h_{j}(X_{j})\cdot 0\big]=0.$
Similarly, for $H_{j}$ and $H_{k\ell}$, conditioning on $X_{j}$ or $(X_{k},X_{\ell})$ and using the defining constraints $\mathbb{E}[h_{k\ell}\mid X_{k}]=\mathbb{E}[h_{k\ell}\mid X_{\ell}]=0$ shows that the inner product vanishes. The same tower-property argument applies to $H_{jk}$ and $H_{k\ell}$, $H_{jk}$ and $H_{j\ell}$, and $H_{jk}$ and $H_{mn}$ for distinct pairs $(j,k)$ and $(m,n)$. Closedness follows from the fact that each subspace is defined by linear constraints expressed in terms of (continuous) conditional expectation operators on $L^{2}$, which preserve closedness. ∎
@@
so $m_{j}(X_{j})$ is centered and depends only on $X_{j}$, hence $m_{j}(X_{j})\in H_{j}$. For $g_{jk}$, note first that
$\mathbb{E}[g_{jk}(X_{j},X_{k})]=\mathbb{E}[m_{jk}(X_{j},X_{k})]-\mathbb{E}[m_{j}(X_{j})]-\mathbb{E}[m_{k}(X_{k})]+\mu=0,$
and
$\mathbb{E}[g_{jk}(X_{j},X_{k})\mid X_{j}]$ $=\mathbb{E}[m_{jk}(X_{j},X_{k})\mid X_{j}]-m_{j}(X_{j})-\mathbb{E}[m_{k}(X_{k})\mid X_{j}]+\mu$
$=\big(\mathbb{E}[f(X)\mid X_{j}]-\mu\big)-m_{j}(X_{j})-\big(\mathbb{E}[f(X)\mid X_{j}]-\mu\big)+\mu$
@@
and symmetrically $\mathbb{E}[g_{jk}(X_{j},X_{k})\mid X_{k}]=0$. Thus $g_{jk}(X_{j},X_{k})\in H_{jk}$, proving (1). ∎
@@
Orthogonality of $r$ to $H_{0}$, the $H_{j}$, and the $H_{jk}$ follows from the tower property of conditional expectation and the defining constraints in (1). For example,
$\mathbb{E}[r(X)]=\mathbb{E}[f(X)]-\mu-\sum_{j}\mathbb{E}[m_{j}(X_{j})]-\sum_{j<k}\mathbb{E}[g_{jk}(X_{j},X_{k})]=0,$
so $r$ is orthogonal to $H_{0}$. For any $h_{j}(X)=h_{j}(X_{j})\in H_{j}$,
$\mathbb{E}[r(X)h_{j}(X_{j})]$ $=\mathbb{E}\big[f(X)h_{j}(X_{j})\big]-\mu\,\mathbb{E}[h_{j}(X_{j})]-\sum_{\ell}\mathbb{E}[m_{\ell}(X_{\ell})h_{j}(X_{j})]$
$\quad-\sum_{k<\ell}\mathbb{E}[g_{k\ell}(X_{k},X_{\ell})h_{j}(X_{j})].$
The second term vanishes since $h_{j}$ is centered. The third term vanishes for $\ell\neq j$ by the same conditioning argument as in Lemma A.3, and for $\ell=j$ by the definition of $m_{j}$ as a centered conditional expectation. The fourth term vanishes by conditioning on $(X_{k},X_{\ell})$ and using the constraints defining $H_{k\ell}$ together with the fact that $h_{j}$ depends only on $X_{j}$. An analogous argument shows orthogonality to each $H_{jk}$, which proves (2). ∎
@@
The decomposition in Proposition A.4 is formally analogous to the standard two-way ANOVA decomposition in designed experiments. In the classical setting, $f$ is a linear model on a finite $2^{d}$ factorial design, and the components in $H_{0}$, $H_{j}$, and $H_{jk}$ correspond to the intercept, main effects, and interaction contrasts of the design matrix. Here, we work in the Hilbert space $L^{2}(P_{X})$ of an arbitrary trained predictor $f$, and the same structure arises from orthogonal projection onto the subspaces of functions depending on at most two coordinates with appropriate centering constraints.
@@ A.3 Discrete ANOVA surrogate as an $L^{2}$ projection
Under Assumptions A.1, A.2, and A.6, the quadratic functional $J$ admits a unique minimizer
@@ A.4 Finite-sample convergence of the surrogate interaction maps
Under Assumptions A.1, A.6, and A.8, the strong law of large numbers implies
@@
Under Assumptions A.1, A.2, A.6, and A.8, the empirical estimator $\widehat{\beta}_{n}$ exists and is unique with probability tending to one as $n\to\infty$, and
@@ A.5 Interaction strength consistency and Top-$K$ / NDCG@K alignment
Suppose Assumptions A.1, A.2, A.6, A.8, A.13, and A.14 hold. Then there exist $L_{0}\in\mathbb{N}$ and $n_{0}\in\mathbb{N}$ such that for all $L\geq L_{0}$ and all $n\geq n_{0}$,
@@ A.7 Further theoretical remarks: grid resolution, misspecification, and higher-order terms
Remark A.20 (Misspecification and dependence on $P_{X}$).
All oracle quantities in Appendix A.2 are defined as orthogonal projections in the Hilbert space $L^{2}(P_{X})$ of the trained predictor $f$. In particular, the main-effect and interaction components $(m_{j},g_{jk})$ and the residual $r$ are defined by their $L^{2}(P_{X})$-best approximation properties, regardless of whether $f$ arises from a ”correctly specified” probabilistic model.
@@
  • The interaction surfaces $g_{jk}$ and strengths $S^{\star}_{jk}$ reflect both the architecture and the data distribution $P_{X}$. If the same predictor $f$ were evaluated under a different input distribution, the oracle decomposition and strengths would, in principle, change.
  • When the true mechanism contains higher-order or non-factorizable structure, these contributions are distributed across the pairwise surfaces $g_{jk}$ and the residual $r$ in the way that minimizes $L^{2}(P_{X})$ error. Our oracle is therefore ”model-centric”: it asks how $f$ can be best approximated by main and pairwise terms under the observed $P_{X}$, rather than asserting that $g_{jk}$ coincides with a causal or structural notion of two-way interaction.
In the presence of misspecification, our evaluation metrics should thus be interpreted as comparing explanation methods against this canonical projection-based oracle, not against a latent ground-truth mechanism that may or may not align with $L^{2}(P_{X})$ geometry.

### v4

@@ 5 Experiments
We compare our ANOVA-based interaction framework, ORACLE, to five SHAP-family (Shapley-based) interaction estimators. ORACLE is not itself a Shapley-based explainer: it does not rely on coalition sampling or game-theoretic axioms, but instead performs an ANOVA-style orthogonal decomposition on a factorial surrogate of the backbone model (Section 3). Nevertheless, ORACLE and the Shapley-based baselines all output a global pairwise interaction score for each feature pair. These scores rely on different definitions, so we do not claim they coincide; instead, we compare them empirically under a common evaluation target, namely the functional-ANOVA oracle in Section 4. ORACLE fits an ANOVA surrogate on a discretized grid and extracts ANOVA interaction maps and interaction strengths as described in Section 3. As SHAP-based baselines, we include the following methods:
  • FSI: the Faithful Shapley interaction index, which provides an axiomatic Shapley-based measure of global interaction strength (Tsai et al., 2023).
  • SHAP-IQ: a unified Monte Carlo scheme for approximating any-order Shapley interactions (Fumagalli et al., 2023).
  • SVARM-IQ: an accelerated variant of SHAP-IQ that uses stratified sampling over subset sizes to reduce variance and computation (Kolpaczki et al., 2024).
  • KS-Int: a KernelSHAP-style weighted least-squares formulation for Shapley interaction indices (Fumagalli et al., 2024).
  • DS-Int: a DeepSHAP-style multi-reference interaction approximation built on gradient-based attributions (Chen et al., 2020).
@@
  • IG@K,B: intervention-oriented utility, defined as the total oracle interaction-strength mass recoverable by selecting up to $B$ pairs from the method’s top-$K$ interactions.
@@ Discussion.
Across three tabular regression benchmarks, ORACLE recovers the backbone’s interaction structure more accurately than SHAP-family interaction estimators, achieving higher NDCG@5 and Xfer-NDCG@5 in nearly all settings along with competitive hotspot localization and scale alignment. Qualitatively, the recovered main effects and interaction maps align with classical DoE-style trends and localized non-additivity patterns (Figures 2 and 3). The intervention metric IG@K,B is more mixed: SHAP methods can match or slightly exceed ORACLE, consistent with their design as marginal-contribution scores tailored to removal-style interventions, whereas ORACLE targets $L^{2}$ fidelity of low-order effect surfaces on the discretized grid. Overall, the results suggest that many tabular predictors admit faithful low-order projections for interaction benchmarking, while the latent image/text case study (Appendix E) highlights reduced competitiveness in highly entangled latent spaces where factorial structure is less interpretable.
@@
ORACLE is grid-based, so performance depends on discretization and can degrade when many features or high resolutions yield sparse cells (our $L$ ablation suggests moderate grids are a practical trade-off). We focus on pairwise effects and use the functional-ANOVA oracle as our evaluation target. Our theoretical analysis assumes independent input coordinates to obtain an orthogonal (Hoeffding-style) decomposition; under strong feature dependence, pairwise projection-based effects can mix with higher-order structure and may deviate from domain- or causal notions of interaction. Finally, our conclusions are drawn from low- to medium-dimensional tabular tasks and MLP backbones, and may not transfer directly to high-dimensional vision or language models with strong local structure or entangled representations.
@@
Scaling beyond low- to medium-dimensional tabular settings will likely require adaptive/sparse grids, dependence-aware bases, and extensions to other backbones and higher-order interactions. On the evaluation side, incorporating robustness under distribution shift, alternative training runs, and architectural changes could yield more comprehensive benchmarks. Finally, hybrid approaches that use interaction maps to guide or regularize SHAP-style estimators may combine structured DoE semantics with the flexibility of Shapley-based indices.
@@ A.1 Additional assumptions and notation
Assumption A.2 (Regular conditional expectations and independent coordinates).
For any subset of coordinates of $X$, regular conditional distributions exist, so that $\mathbb{E}[f(X)\mid X_{j}=x_{j}]$ and $\mathbb{E}[f(X)\mid X_{j}=x_{j},X_{k}=x_{k}]$ are well-defined for $P_{X_{j}}$- and $P_{X_{j},X_{k}}$-almost every $(x_{j},x_{k})$. Moreover, the coordinates $(X_{1},\dots,X_{d})$ are mutually independent under $P_{X}$ (equivalently, $P_{X}=\bigotimes_{j=1}^{d}P_{X_{j}}$).
@@
Under Assumptions A.1A.3 and A.2, the oracle main and pairwise components in the main text can be written as
@@ A.2 Oracle ANOVA decomposition: existence and representation
Orthogonality of $H_{0}$ and each $H_{j},H_{jk}$ follows from the centering constraints. For $H_{j}$ and $H_{\ell}$ with $j\neq\ell$, take $h_{j}(X_{j})\in H_{j}$ and $h_{\ell}(X_{\ell})\in H_{\ell}$. By independence (Assumption A.2), $\mathbb{E}[h_{\ell}(X_{\ell})\mid X_{j}]=\mathbb{E}[h_{\ell}(X_{\ell})]=0$, hence
$\langle h_{j}(X_{j}),h_{\ell}(X_{\ell})\rangle=\mathbb{E}\!\left[h_{j}(X_{j})\,\mathbb{E}[h_{\ell}(X_{\ell})\mid X_{j}]\right]=0.$
Next, consider $H_{j}$ and $H_{k\ell}$. If $j\notin\{k,\ell\}$, independence implies $\mathbb{E}[h_{k\ell}(X_{k},X_{\ell})\mid X_{j}]=\mathbb{E}[h_{k\ell}(X_{k},X_{\ell})]=0$ (since $h_{k\ell}$ is centered), so the inner product is zero. If $j\in\{k,\ell\}$, the defining conditional-centering constraints of $H_{k\ell}$ (e.g. $\mathbb{E}[h_{k\ell}\mid X_{k}]=0$) yield orthogonality by the tower property. A similar argument applies to $H_{jk}$ and $H_{mn}$: when the index pairs are disjoint, independence reduces the relevant conditional expectation to the (centered) marginal mean; when they overlap, the conditional-centering constraints enforce orthogonality.
Finally, $\mathcal{H}_{\leq 2}$ is closed because it is finite-dimensional (in the discretized case) or, in the population $L^{2}$ setting, it is defined by linear constraints given by conditional-expectation operators that are continuous on $L^{2}$. ∎
@@
so $m_{j}(X_{j})$ is centered and depends only on $X_{j}$, hence $m_{j}(X_{j})\in H_{j}$.
For $g_{jk}$, we first note $\mathbb{E}[g_{jk}(X_{j},X_{k})]=0$ by construction. Moreover, by the tower property,
$\mathbb{E}\!\left[m_{jk}(X_{j},X_{k})\mid X_{j}\right]=\mathbb{E}\!\left[\mathbb{E}[f(X)\mid X_{j},X_{k}]-\mu\,\Big|\,X_{j}\right]=\mathbb{E}[f(X)\mid X_{j}]-\mu=m_{j}(X_{j}),$
and similarly $\mathbb{E}[m_{jk}(X_{j},X_{k})\mid X_{k}]=m_{k}(X_{k})$. Using Assumption A.2, we have $X_{k}\perp X_{j}$ and hence $\mathbb{E}[m_{k}(X_{k})\mid X_{j}]=\mathbb{E}[m_{k}(X_{k})]=0$. Therefore
$\mathbb{E}[g_{jk}(X_{j},X_{k})\mid X_{j}]$ $=\mathbb{E}[m_{jk}(X_{j},X_{k})\mid X_{j}]-m_{j}(X_{j})-\mathbb{E}[m_{k}(X_{k})\mid X_{j}]+\mu-\mu$
$=m_{j}(X_{j})-m_{j}(X_{j})-0$
@@
and symmetrically $\mathbb{E}[g_{jk}(X_{j},X_{k})\mid X_{k}]=0$. Thus $g_{jk}(X_{j},X_{k})\in H_{jk}$. ∎
@@
Orthogonality of $r$ to $H_{0}$ follows from
$\mathbb{E}[r(X)]=\mathbb{E}[f(X)]-\mu-\sum_{j}\mathbb{E}[m_{j}(X_{j})]-\sum_{j<k}\mathbb{E}[g_{jk}(X_{j},X_{k})]=0.$
Let $h_{j}(X)=h_{j}(X_{j})\in H_{j}$. Using linearity,
$\mathbb{E}[r(X)\,h_{j}(X_{j})]=\mathbb{E}[f(X)\,h_{j}(X_{j})]-\mu\,\mathbb{E}[h_{j}(X_{j})]-\sum_{\ell}\mathbb{E}[m_{\ell}(X_{\ell})\,h_{j}(X_{j})]-\sum_{k<\ell}\mathbb{E}[g_{k\ell}(X_{k},X_{\ell})\,h_{j}(X_{j})].$
The second term vanishes since $\mathbb{E}[h_{j}(X_{j})]=0$.
For $\ell\neq j$, independence gives $X_{\ell}\perp X_{j}$, hence $\mathbb{E}[m_{\ell}(X_{\ell})\,h_{j}(X_{j})]=\mathbb{E}[m_{\ell}(X_{\ell})]\,\mathbb{E}[h_{j}(X_{j})]=0$. For $\ell=j$,
$\mathbb{E}[f(X)\,h_{j}(X_{j})]=\mathbb{E}\!\left[\mathbb{E}[f(X)\mid X_{j}]\,h_{j}(X_{j})\right]=\mathbb{E}\!\left[(m_{j}(X_{j})+\mu)\,h_{j}(X_{j})\right]=\mathbb{E}[m_{j}(X_{j})\,h_{j}(X_{j})],$
since $\mathbb{E}[h_{j}(X_{j})]=0$, which cancels the $\ell=j$ contribution.
For the interaction terms: if $j\notin\{k,\ell\}$, then $(X_{k},X_{\ell})\perp X_{j}$ and thus $\mathbb{E}[g_{k\ell}(X_{k},X_{\ell})\,h_{j}(X_{j})]=0$ by factorization. If $j=k$ (or $j=\ell$), then conditioning on $X_{j}$ and using $\mathbb{E}[g_{j\ell}(X_{j},X_{\ell})\mid X_{j}]=0$ yields
$\mathbb{E}[g_{j\ell}(X_{j},X_{\ell})\,h_{j}(X_{j})]=\mathbb{E}\!\left[h_{j}(X_{j})\,\mathbb{E}[g_{j\ell}(X_{j},X_{\ell})\mid X_{j}]\right]=0.$
Therefore $\mathbb{E}[r(X)\,h_{j}(X_{j})]=0$ for all $h_{j}\in H_{j}$.
An analogous argument shows orthogonality to each $H_{jk}$: for any $h_{jk}(X_{j},X_{k})\in H_{jk}$, expanding $\mathbb{E}[r\,h_{jk}]$ yields only terms that vanish either by independence (disjoint index sets) or by the conditional-centering constraints defining $H_{jk}$ (shared indices). Hence $r\in R$. ∎
@@
The decomposition in Proposition A.5 is formally analogous to the standard two-way ANOVA decomposition in designed experiments. In the classical setting, $f$ is a linear model on a finite $2^{d}$ factorial design, and the components in $H_{0}$, $H_{j}$, and $H_{jk}$ correspond to the intercept, main effects, and interaction contrasts of the design matrix. Here, we work in the Hilbert space $L^{2}(P_{X})$ of an arbitrary trained predictor $f$, and the same structure arises from orthogonal projection onto the subspaces of functions depending on at most two coordinates with appropriate centering constraints (under the product measure in Assumption A.2).
@@ A.3 Discrete ANOVA surrogate as an $L^{2}$ projection
Under Assumptions A.1, A.3, and A.7, the quadratic functional $J$ admits a unique minimizer
@@ A.4 Finite-sample convergence of the surrogate interaction maps
Under Assumptions A.1, A.7, and A.9, the strong law of large numbers implies
@@
Under Assumptions A.1, A.3, A.7, and A.9, the empirical estimator $\widehat{\beta}_{n}$ exists and is unique with probability tending to one as $n\to\infty$, and
@@ A.5 Interaction strength consistency and Top-$K$ / NDCG@K alignment
Suppose Assumptions A.1, A.3, A.7, A.9, A.14, and A.15 hold. Then there exist $L_{0}\in\mathbb{N}$ and $n_{0}\in\mathbb{N}$ such that for all $L\geq L_{0}$ and all $n\geq n_{0}$,
@@ A.7 Further theoretical remarks: grid resolution, misspecification, and higher-order terms
Remark A.21 (Misspecification and dependence on the reference measure).
All oracle quantities in Appendix A.2 are defined as orthogonal projections in the Hilbert space $L^{2}(P_{X})$ of the trained predictor $f$. In this appendix, the reference measure $P_{X}$ is the product measure specified in Assumption A.2.
@@
  • The interaction surfaces $g_{jk}$ and strengths $S^{\star}_{jk}$ reflect both the architecture/predictor $f$ and the chosen reference distribution. If the same predictor $f$ were evaluated under a different set of marginals $\{P_{X_{j}}\}$ (and hence a different product measure), the oracle decomposition and strengths would, in principle, change.
  • When the true mechanism contains higher-order or non-factorizable structure, these contributions are distributed across the pairwise surfaces $g_{jk}$ and the residual $r$ in the way that minimizes $L^{2}(P_{X})$ error under the product measure. Our oracle is therefore ”model-centric” and ”reference-measure-centric”: it asks how $f$ can be best approximated by main and pairwise terms under the product reference distribution, rather than asserting that $g_{jk}$ coincides with a causal or structural notion of two-way interaction.
Accordingly, our evaluation metrics should be interpreted as comparing explanation methods against this canonical projection-based oracle under the product measure.